OpenStudy (anonymous):

rewrite: (3x)^5=2y as a logarithm

OpenStudy (anonymous):

first distribute the power into the brackets

OpenStudy (anonymous):

so like 3^5=243 243x^5=2y if you need to set = to 0 then subtract 2y 243x^5-2y=0

OpenStudy (anonymous):

as a logarithm?

OpenStudy (anonymous):

no, I completely read over the word logarithm so that is wrong